6x2 - 20 + 7x = 0

Answer by mananth(16946)

6x2 - 20 + 7x = 0
6x2 +7x -20 = 0
6x^2+15x-8x-20
3x(2x+5)-4(2x+5)=0
(2x+5)(3x-4)=0
x=-5/2 , or x=4/3
...
m.ananth@hotmail.ca